OpenStudy (anonymous):

Use the Subtraction Property of Equality to complete the following statement: If 10x +6 = 21, then ____ = 15 10x 1.5 10x – 6 10x + 6

OpenStudy (mayankdevnani):

@Margaret17 first find the value of x \[\huge 10x+6=21\]

OpenStudy (anonymous):

the value of x is 1.5 for that equation

OpenStudy (mayankdevnani):

yeah!! correct

OpenStudy (mayankdevnani):

then take this value and then put this value in options .If they satisfy the question's value, then that option is correct

OpenStudy (anonymous):

okay

OpenStudy (mayankdevnani):

ok..so what is the answer??

OpenStudy (mayankdevnani):

@Margaret17

OpenStudy (anonymous):

1.5 ??

OpenStudy (mayankdevnani):

first put this value in option A \[\huge 10x=10 \times 1.5\] so what is the answer??

OpenStudy (anonymous):

1.5

OpenStudy (mayankdevnani):

wrong!!! what will you get when you multiply this \[\huge 10 \times 1.5=???\]

OpenStudy (anonymous):

15

OpenStudy (mayankdevnani):

correct..so your answer is option A